OpenStudy (anonymous):

g(t)=(t-4)ln(t+5) find g'(t) & g"(t)

OpenStudy (anonymous):

i thought i understood this problem but no... i thought ln(t+5) turns to (1/(t+5)) and (t-4) turns to 1 so the answer would be 1/(t+5) for g'(t) but its not...

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

sorry had to tag you lol

OpenStudy (ash2326):

we have \[g(t)=t-4 \ln(t-5)\] using product rule we get \[g'(t)=(t-4)\frac{1}{t-5}+ \ln (t-5)\]

OpenStudy (ash2326):

You try for g''(t)

OpenStudy (anonymous):

would it be 1/(t+5) + 1/(t+5) so 2/(t+5)??

OpenStudy (ash2326):

No, for the first term use division rule (u/v)'= (vu'-uv')/v^2 \[t-4/t-5= ( (t-5)(1)-(t-4)(1))/(t-5)^2

OpenStudy (anonymous):

im not seeing it. can u fill in the steps u skipped.

OpenStudy (anonymous):

sorry if im being complicated

OpenStudy (ash2326):

We have \[g'(t)=\frac{t-4}{t-5}+\ln (t-5)\] let's find g''(t) \[g''(t)=\frac{(t-5)\frac{d}{dt} (t-4)-(t-4)\frac{d}{dt}(t-5)}{(t-5)^2}+\frac{1}{t-5}\] we get \[g''(t)=\frac{(t-5)\times 1-(t-4) \times 1}{(t-5)^2}+\frac{1}{t-5}\] \[g''(t)=\frac{t-5-t+4}{(t-5)^2}+\frac{1}{t-5}\] we get \[g''(t)=\frac{-1}{(t-5)^2}+\frac{1}{t-5}\]

OpenStudy (anonymous):

i thought it was PLUS 5 (t+5)

OpenStudy (ash2326):

@mariomintchev did you understand?

OpenStudy (anonymous):

why is it t-5 not t+5?

OpenStudy (ash2326):

Oh I didn't see that it's t+5 . Sorry replace (t-5) by (t+5) everywhere

OpenStudy (anonymous):

so 9/(t+5)^2 + 1/(t+5) is the final answer

OpenStudy (ash2326):

yeah :D

OpenStudy (anonymous):

my hw website isnt accepting that answer

OpenStudy (ash2326):

It can also be written as \[g''(t)= \frac{9}{(t+5)^2}+\frac{1}{t+5}=\frac{9}{(t+5)^2}+\frac{t+5}{(t+5)^2}\] or \[g''(t)= \frac{t+14}{(t+5)^2}\] try this

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

any other ways i could put it in?

OpenStudy (anonymous):

@ash2326 you online?

OpenStudy (anonymous):

\[g(t)=(t-4)ln(t+5)\] \[g(x)=t\ln(t+5)-4\ln(t+5)\] to start, then \[g'(t)=\ln(t+5)+\frac{t}{t+5}-\frac{4}{t+5}\] or if you prefer \[g'(x)=\ln(t+5)+\frac{t-4}{t+5}\] n replete \[g''(x)=\frac{1}{t+5}+\frac{9}{(t+5)^2}\] or if you want to add the fractions \[\frac{t+4+9}{(t+5)^2}\]

OpenStudy (anonymous):

actually this is exactly what ash wrote above

OpenStudy (anonymous):

@amistre64 this is the final one that we've been struggling with

OpenStudy (anonymous):

any ideas?

OpenStudy (saifoo.khan):

Sorry mario. Im not good with calc. :(

OpenStudy (anonymous):

:(

OpenStudy (amistre64):

didja get it yet?

OpenStudy (anonymous):

nooooooooooooo

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

everyones getting similar answers but my hw site isnt accepting anything i put in. ive put in over 20 answers. ha

OpenStudy (amistre64):

this one aint asking for values

OpenStudy (amistre64):

i can redo whats already been done; but that wont really change things will it

OpenStudy (amistre64):

unless you gave the wrong problem, theres not much more to do

OpenStudy (anonymous):

g(t)=(t+4)ln(t+5)

OpenStudy (anonymous):

g'(t) is ln(t+5)+(t+4)/(t+5)

OpenStudy (anonymous):

i need g''(t)

OpenStudy (amistre64):

then do it again :) nothings stopping you from applying derivative again to it

OpenStudy (amistre64):

show me what you get and lets see where you having the issues at

OpenStudy (anonymous):

everywhere

OpenStudy (anonymous):

been on this problem forever

OpenStudy (amistre64):

yes; everywhere

OpenStudy (anonymous):

Ok I got: \[g'(t)=\ln{(t+5)}+\frac{t-4}{t+5}\]

OpenStudy (amistre64):

\[g'(t) = ln(t+5)+\frac{(t+4)}{(t+5)}\] derive it thru again for g''

OpenStudy (amistre64):

lets use the last edition of the problem this time tho ...

OpenStudy (anonymous):

and \[g''(t)=\frac{t+14}{(t+5)^2}\]

OpenStudy (amistre64):

the t-4 is from the first post

OpenStudy (anonymous):

is not t-4?

OpenStudy (anonymous):

i end up with 9/(t+5)^2 + 1/(t+5)

OpenStudy (anonymous):

what is the function you want to derivate?

OpenStudy (anonymous):

its PLUS 4. SORRY!

OpenStudy (anonymous):

well in that case I got \[g''(t)=\frac{t+2}{(t+5)^2}\]

OpenStudy (anonymous):

and \[g'(t)=\ln{(t+5)+\frac{t+4}{t+5}\]

OpenStudy (amistre64):

lets see what I get :) \[g'(t) = ln(t+5)+\frac{(t+4)}{(t+5)}\] \[g''(t) = ln(t+5)'+[(t+4)(t+5)^{-1}]'\] \[g''(t) = (t+5)^{-1}+(t+4)'(t+5)^{-1}+(t+4)(t+5)' ^{-1}\] \[g''(t) = (t+5)^{-1}+t(t+5)^{-1}-(t+4)(t+5)^{-2}\] \[g''(t) = ((t+5)+t(t+5)-(t+4))(t+5)^{-2}\] \[g''(t) = \frac{(t+5)+t(t+5)-(t+4)}{(t+5)^{2}}\] \[g''(t) = \frac{t+5+tt+5t-t-4}{(t+5)^{2}}\] \[g''(t) = \frac{t^2 +5t+1}{(t+5)^{2}}\]

OpenStudy (anonymous):

\[(t+4)'\neq t\]\@amistre64

OpenStudy (amistre64):

yeah, was just getting to that :)

OpenStudy (amistre64):

t+6 up top

OpenStudy (amistre64):

wonderful piece of work ruint by a lousy t :)

OpenStudy (anonymous):

yeah a lot of latex work =/

OpenStudy (anonymous):

but what is the answer you get then?

OpenStudy (ash2326):

I think it'd be \[g''(t)=\frac{1}{t+5}+\frac{1}{(t+5)^2}\] we get \[g''(t)=\frac{(t+6)}{(t+5)^2}\]

OpenStudy (amistre64):

10-4 = 6 soo \[g''=\frac{t+6}{(t+5)^2}\]

OpenStudy (anonymous):

i love you all!!!

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

Yeah, you're awesome!